\small

\begin{thebibliography}{99}
\bibitem{DHST11}
Denton, T., Hivert, F., Schilling, A., Thi\'ery, N.M.: On the representation theory of finite $\mathcal{J}$-trivial monoids. S\'eminaire Lotharingien de Combinatoire \textbf{64}, article no. B64d (2011)

\bibitem{GaMa09}
Ganyushkin, O., Mazorchuk, V.: Classical Finite Transformation Semigroups: An Introduction. Springer, London (2009)

\bibitem{GaoZhangLuo22}
Gao, M., Zhang, W.T., Luo, Y.F.: Finite basis problem for Catalan monoids with involution. Int. J. Algebra Comput. 32(06), 1161-–1177 (2022)

\bibitem{GSV25}
Gusev, S.V., Sapir, O.B., Volkov, M.V.:   Strongly nonfinitely based monoids, J. Combinatorial Algebra, \textbf{9}(1-2), 129--144 (2025)

\bibitem{Hi93}
Higgins, P.M.: {Combinatorial results for semigroups of order-preserving mappings}. Math. Proc. Cambridge Phil. Soc. \textbf{113}(2), 281--296 (1993)

\bibitem{HiTh09}
Hivert, F., Thi\'ery, N.M.: The Hecke group algebra of a Coxeter group and its representation theory. J. Algebra \textbf{321}, 2230--2258 (2009)

\bibitem{JF19}
Johnson, M., Fenner, P.: Identities in unitriangular and gossip monoids. Semigroup Forum \textbf{98}, 338--354 (2019)

\bibitem{KP10}
Kl\'ima, O., Pol\'ak, L.: Hierarchies of piecewise testable languages. Int. J. Found. Comput. Sci. \textbf{21}(4), 517--533 (2010)

\bibitem{LuoJinZhang25}
Luo, Y.F., Jin, Z.F., Zhang, W.T.: Finite Gr\"obner--Shirshov bases and faithful representations for Catalan monoids. Commun. Algebra \textbf{54}(1), 185--195 (2026) 

\bibitem{SaVo23}
Sapir, O.B., Volkov, M.V.: Catalan monoids inherently nonfinitely based relative to finite $R$-trivial semigroups. {J. Algebra} \textbf{633}, 138--171 (2023)

\bibitem{So96}
Solomon, A.: {Catalan monoids, monoids of local endomorphisms, and their presentations}. Semigroup Forum \textbf{53}(3),  351--368 (1996)

\bibitem{Stanley}
Stanley, R.P.: Catalan Numbers. Cambridge University Press, Cambridge (2015)

\bibitem{Ts90}
Tsaranov, S.V.: Representation and classification of Coxeter monoids. Eur. J. Comb. \textbf{11}(2), 189--204 (1990)

\bibitem{Vo04}
Volkov, M.V.: Reflexive relations, extensive transformations and piecewise testable languages of a given height.  Int. J. Algebra Comput. \textbf{14} (5-6),  817--827 (2004)

\bibitem{Vo25}
Volkov, M.V.: Identities of triangular Boolean matrices. Preprint arXiv:2412.16113v4 (2025)
\end{thebibliography}

\documentclass[11pt,reqno]{amsart}
\usepackage[foot]{amsaddr}
\usepackage{amssymb,amsmath,amsfonts,mathptmx,cite,mathrsfs,graphicx,rotating,url}
\usepackage{mathdots}

\newtheorem{theorem}{Theorem}
\newtheorem{cor}[theorem]{Corollary}

\theoremstyle{definition}

\newtheorem*{remark}{Remark}

\usepackage{tikz}
\usetikzlibrary{arrows}
\usetikzlibrary{arrows,shapes}

\DeclareSymbolFont{rsfscript}{OMS}{rsfs}{m}{n}
\DeclareSymbolFontAlphabet{\mathrsfs}{rsfscript}

\newcommand{\ais}{ai-semi\-ring}

\newcommand{\mC}{\mathcal{C}}

\newcommand{\mS}{\mathcal{S}}
\newcommand{\mT}{\mathcal{T}}

\makeatletter

\renewcommand*\subjclass[1]{\def\@subjclass{#1}{\@xp\let\@xp\subjclassname\csname subjclassname@#1\endcsname}}

\renewcommand{\subjclassname}{\textup{2020} Mathematics Subject Classification}

\makeatother

\title[A matrix representation for Catalan semirings]{A new Boolean matrix representation\\ for Catalan semirings}
\author{Mikhail Volkov}

\address{{\normalfont 620075 Ekaterinburg, Russia}}
\email{m.v.volkov@urfu.ru}

\date{}

\thanks{Supported by the Ministry of Science and Higher Education of the Russian Federation, project FEUZ-2023-2022}

\subjclass{16Y60, 15B34, 06F05, 20M20}
\keywords{Catalan monoid, Catalan semiring, Boolean matrix representation}
\begin{document}

\begin{abstract}
We construct a faithful representation of the semiring of all order-preserving decreasing transformations of a chain with $n+1$ elements by Boolean upper triangular $n\times n$-matrices.
\end{abstract}

\maketitle

\subsection*{Catalan monoids} Let $[n]$ stand for the set $\{1,2,\dots,n\}$ of the first $n$ positive integers. We consider the set $[n]$ with its usual order: $1<2<\dots<n$. A transformation $\alpha\colon[n]\to[n]$ is \emph{order-preserving} if $i\le j$ implies $i\alpha\le j\alpha$ for all $i,j\in[n]$ and \emph{extensive} if $i\le i\alpha$ for every $i\in[n]$. Clearly, if two transformations have either of the properties of being order-preserving or extensive, then so does their product, and the identity transformation has both properties. Hence, the set $C_n$ of all extensive order-preserving transformations of $[n]$ forms a submonoid in the monoid of all transformations of $[n]$. This submonoid often appears in the literature as the \emph{Catalan} monoid. The name is justified by the cardinality of $C_n$ being the $n$-th Catalan number $\frac1{n+1}\binom{2n}n$; see \cite[Theorem 14.2.8(i)]{GaMa09}.

A transformation $\alpha\colon[n]\to[n]$ is called \emph{decreasing} (or \emph{parking}) if $i\alpha\le i$ for all $i\in[n]$. The set $C^-_n$ of all decreasing order-preserving transformations of $[n]$ also forms a submonoid in the monoid of all transformations of $[n]$. The submonoid $C^-_n$ differs from the submonoid $C_n$ if $n>1$, but the two monoids are isomorphic because $C^-_n$ is nothing more than the monoid of all extensive order-preserving transformations of the set $\{1,2,\dots,n\}$ equipped with the `opposite' order $1>2>\dots>n$. Slightly abusing terminology, we collectively refer to monoids in both series $\{C_n\}_{n=1,2,\dotsc}$ and $\{C^-_n\}_{n=1,2,\dotsc}$ as Catalan monoids.

Catalan monoids have been intensively studied from various viewpoint (and under various names); we mention \cite{DHST11,GaoZhangLuo22,GSV25,Hi93,HiTh09,LuoJinZhang25,SaVo23,So96,Ts90,Vo04} as samples of such studies.

\subsection*{Catalan semirings} An \emph{additively idempotent semiring} (\ais, for short) is an algebra $(A,+,\cdot)$ with binary addition $+$ and multiplication $\cdot$ such that the additive reduct $(A,+)$ is a commutative and idempotent semigroup, the multiplicative reduct $(A,\cdot)$ is a semigroup, and multiplication distributes over addition on the left and right.

The set $O_n$ of all order-preserving transformations of $[n]$ is a submonoid in the monoid of all transformations of $[n]$; clearly, both $C_n$ and $C^-_n$ are submonoids of the monoid $O_n$. We define, for all $\alpha,\beta\in O_n$ and $i\in[n]$,
\[
i(\alpha+\beta):=\max\{i\alpha,i\beta\}.
\]
(Here and below the sign $:=$ stands for equality by definition; thus, $A:=B$ means that $A$ is defined as $B$.) It is well known (and easy to verify) that $\alpha+\beta\in O_n$. Equipped with this addition, $O_n$ becomes an \ais{} in which both $C_n$ and $C^-_n$ form subsemirings. Hence, Catalan monoids admit a natural \ais{} structure. We denote the \ais{}s defined on $C_n$ and $C^-_n$ by $\mC_n$ and $\mC^-_n$, respectively, and refer to them as \emph{Catalan} semirings. Notice that while for each $n$, the monoids $C_n$ and $C^-_n$ are isomorphic, the semirings $\mC_n$ and $\mC^-_n$ are not isomorphic if $n>1$, and moreover, their additive reducts are not isomorphic if $n>2$.

For $\alpha,\beta\in O_n$, let
\begin{equation}\label{eq:semilattice order}
\alpha\leqslant \beta\iff \alpha+\beta=\beta\iff i\alpha\le i\beta\ \text{ for all }\  i\in[n].
\end{equation}
The relation $\leqslant$ is a partial order on the set $O_n$, under which $O_n$ becomes a distributive lattice. The addition on $O_n$ can be recovered from the order \eqref{eq:semilattice order}---for all $\alpha,\beta\in O_n$, their sum $\alpha+\beta$ coincides with the least upper bound of $\alpha$ and $\beta$ with respect to $\leqslant$.

For illustration, the next picture shows the Hasse diagram of the lattice $(O_3,\leqslant)$. Each transformation $\alpha\in O_3$ is encoded by the triple $1\alpha 2\alpha 3\alpha$.
\begin{center}
\begin{tikzpicture}[x=1.6cm,y=1.0cm, every node/.style={font=\small}]
  \node (111) at (0,0) {$111$};

  \node (112) at (0,1) {$112$};

  \node (113) at (-1,2) {$113$};
  \node (122) at (1,2) {$122$};

  \node (123) at (0,3) {$123$};
  \node (222) at (2,3) {$222$};

  \node (133) at (-1,4) {$133$};
  \node (223) at (1,4) {$223$};

  \node (233) at (0,5) {$233$};

  \node (333) at (0,6) {$333$};

  \draw (111) -- (112);

  \draw (112) -- (113);
  \draw (112) -- (122);

  \draw (113) -- (123);
  \draw (122) -- (123);
  \draw (122) -- (222);

  \draw (123) -- (133);
  \draw (123) -- (223);
  \draw (222) -- (223);

  \draw (133) -- (233);
  \draw (223) -- (233);

  \draw (233) -- (333);
\end{tikzpicture}
\end{center}
In particular, the identity transformation $\varepsilon$ is encoded by the triple 123. The five transformations $\alpha$ with $\alpha\geqslant\varepsilon$ form the semiring $\mC_3$, while the five transformations $\beta$ with $\beta\leqslant\varepsilon$ constitute the semiring $\mC^-_3$. The general picture looks the same: the transformations $\alpha\in O_n$ with $\alpha\geqslant\varepsilon$ form the semiring $\mC_n$, while the transformations $\beta\in O_n$ with $\beta\leqslant\varepsilon$ constitute the semiring $\mC^-_n$.

\subsection*{Semirings of Boolean matrices} A \emph{Boolean} matrix is a square matrix with entries $0$ and~$1$ only. The addition and multiplication of such matrices are as usual, except that addition and multiplication of the entries are defined as $x+y:=\max\{x,y\}$ and $x\cdot y:=\min\{x,y\}$. For each $n$, the set of all Boolean $n\times n$-matrices forms an \ais{} under matrix addition and multiplication.

A Boolean matrix $\bigl(a_{ij}\bigr)_{n\times n}$ is called \emph{upper triangular} if $a_{ij}=0$ for all $1\le j<i\le n$. The set of all upper triangular Boolean $n\times n$-matrices is closed under matrix addition and multiplication so it forms a subsemiring of the \ais{} of all Boolean $n\times n$-matrices. We denote by $T_n$ and $\mT_n$ the monoid and, respectively, the \ais{} of all upper triangular Boolean $n\times n$-matrices.

By a \emph{faithful representation} of a semigroup $(A,\cdot)$, respectively, of a \ais{} $(A,+,\cdot)$ in the monoid $T_n$, respectively, in the \ais{} $\mT_n$, we mean an injective map $A\to T_n$ which is a semigroup, respectively, semiring homomorphism.

\subsection*{A faithful representation $\mC_n\to\mT_n$} Any transformation $\alpha\colon[n]\to[n]$ can be represented as the Boolean $n\times n$-matrix $B(\alpha)$ whose entry in position $(i,j)$ is 1 if $i\alpha=j$ and 0 otherwise. The map $\alpha\mapsto B(\alpha)$ is an injective homomorphism from the monoid of all transformations of $[n]$ in the monoid of all Boolean $n\times n$-matrices. If $\alpha$ is extensive, then $i\le i\alpha$ whence the matrix $B(\alpha)$ is upper triangular. Therefore, the restriction of $\alpha\mapsto B(\alpha)$ to the monoid $C_n$ is a faithful representation of $C_n$ in the monoid $T_n$ of all upper triangular Boolean $n\times n$-matrices. However, the map $\alpha\mapsto B(\alpha)$ fails to respect addition so that it is not a representation of the semiring $\mC_n$ in the semiring $\mT_n$.

Ond\v{r}ej Kl\'ima and Libor Pol\'ak (see \cite[Section~5]{KP10}) found another faithful representation of the monoid $C_n$ in the monoid $T_n$. (The same representation was rediscovered by Marianne Johnson and Peter Fenner \cite[Lemma 5.1]{JF19} and again by Yan Feng Luo, Zhen Feng Jin, and Wen Ting Zhang \cite[Theorem 5.4]{LuoJinZhang25}.) Namely, let $S_n$ stand for the set of all \emph{stair triangular} matrices, that is, upper triangular Boolean matrices $\bigl(a_{ij}\bigr)_{n\times n}$ satisfying: $a_{ii}=1$ for all $i=1,\dots,n$, and if $a_{ij} = 1$ for some $1\le i < j\le n$, then
\[
a_{ii+1}=\cdots=a_{ij}=a_{i+1j}=\cdots=a_{j-1j}=1.
\]
The set $S_n$ is closed under the usual multiplication of Boolean matrices and contains the identity matrix. Thus, $S_n$ forms a monoid and the map $\alpha\mapsto S(\alpha)$, where the $(i,j)$-th entry of the matrix $S(\alpha)$ equals 1 if and only if $i\le j\le i\alpha$, is an isomorphism between the monoids $C_n$ and $S_n$.

It is easy to see that the set $S_n$ is also closed under the entry-wise addition of Boolean matrices, so it also forms a subsemiring $\mS_n$ in the semiring $\mT_n$. Although this was not explicitly mentioned in \cite{KP10}, the above isomorphism  $\alpha\mapsto S(\alpha)$ between the monoids $C_n$ and $S_n$ is, in fact, an isomorphism between the \ais{}s $\mC_n$ and $\mS_n$, hence, a faithful representation of the semiring $\mC_n$ in the semiring $\mT_n$.

\subsection*{A faithful representation $\mC^-_{n+1}\to\mT_n$} A Boolean matrix $\bigl(a_{ij}\bigr)_{n\times n}$ is called \emph{lower triangular} if $a_{ij}=0$ for all $1\le i<j\le n$. The set of all lower triangular Boolean $n\times n$-matrices also forms a subsemiring of the \ais{} of all Boolean $n\times n$-matrices; we denote this subsemiring by $\mT^-_n$. If $P$ is the $n\times n$-matrix with 1s on the antidiagonal and 0s elsewhere,
\[
P:=\begin{pmatrix}
0 & 0 & \cdots & 0 & 1\\
0 & 0 & \cdots & 1 & 0\\
\vdots & \vdots & \iddots & \vdots & \vdots\\
0 & 1 & \cdots & 0 & 0\\
1 & 0 & \cdots & 0 & 0\\
\end{pmatrix},
\]
then it is known (and easy to verify) that the map $M\mapsto PMP$ is an isomorphism between the \ais{}s $\mT^-_n$ and $\mT_n$. Therefore, to obtain a faithful representation of the semiring $\mC^-_{n+1}$ in $\mT_n$, it suffices to give a faithful representation in $\mT^-_n$, which is notationally easier.

To each transformation $\alpha\in\mC^-_{n+1}$, we assign a Boolean $n\times n$-matrix $M(\alpha)$ whose $i$-th row consists of $(i+1)\alpha-1$ left-justified 1s; in other words, the $i$-th row is the unary representation of the number $(i+1)\alpha-1$, left-justified. The rule can also be expressed by saying that the $(i,j)$-th entry of the matrix $M(\alpha)$ equals 1 if and only if $(i+1)\alpha-1\ge j$.

For illustration, the next table shows the map $\alpha\mapsto M(\alpha)$ for transformations in $\mC^-_3$.
\[
\begin{array}{c|c|c|c|c|c|}
\text{Transformation}\ \alpha\in\mC^-_3&&&&&\\[-.5em]
\raisebox{5pt}{(encoded as \  $1\alpha 2\alpha 3\alpha$)} & \raisebox{10pt}{111} & \raisebox{10pt}{112} & \raisebox{10pt}{113} & \raisebox{10pt}{122} & \raisebox{10pt}{123} \\[-1em]
& \downarrow & \downarrow & \downarrow & \downarrow & \downarrow \\[-1em]
& & & & & \\
\text{Boolean matrix } M(\alpha)&
\begin{pmatrix}0 & 0 \\ 0 & 0\end{pmatrix} &
\begin{pmatrix}0 & 0 \\ 1 & 0\end{pmatrix} &
\begin{pmatrix}0 & 0 \\ 1 & 1\end{pmatrix} &
\begin{pmatrix}1 & 0 \\ 1 & 0\end{pmatrix} &
\begin{pmatrix}1 & 0 \\ 1 & 1\end{pmatrix}
\end{array}
\]

\begin{theorem}
\label{thm:construction}
The map $\alpha\mapsto M(\alpha)$ is a faithful representation of the semiring $\mC^-_{n+1}$ in the semiring $\mT^-_n$.
\end{theorem}

\begin{proof}
As $\alpha\in\mC^-_{n+1}$ is a decreasing transformation, $(i+1)\alpha\le i+1$ for each $i=1,\dots,n$. Hence,
$(i+1)\alpha-1\le i$ and one must have $j\le i$ for the condition $(i+1)\alpha-1\ge j$ to hold. Thus, only the $i$ first entries in the $i$-th row of the matrix $M(\alpha)$ may be non-zero, and therefore, $M(\alpha)$ is a lower triangular matrix.

It is clear that the map $\alpha\mapsto M(\alpha)$ is one-to-one and preserves addition. To verify that $M(\alpha\beta)=M(\alpha)M(\beta)$ for all transformations $\alpha,\beta\in\mC^-_{n+1}$, let $M(\alpha):=\bigl(a_{ij}\bigr)_{n\times n}$, $M(\beta):=\bigl(b_{ij}\bigr)_{n\times n}$, $M(\alpha\beta):=\bigl(c_{ij}\bigr)_{n\times n}$, and $M(\alpha)M(\beta):=\bigl(d_{ij}\bigr)_{n\times n}$. For all $i,j\in\{1,\dots,n\}$,
\begin{align}
 \notag d_{ij}=\sum_k^{n} a_{ik}b_{kj}=1&\iff \exists k\ (a_{ik}=1\ \& \ b_{kj}=1)\\
 \notag &\iff \exists k\  \bigl((i+1)\alpha-1\ge k\ \& \  (k+1)\beta-1\ge j\bigr)\\
 \label{eq:statement}&\iff \exists k\  \bigl((i+1)\alpha\ge k+1\ \& \  (k+1)\beta\ge j+1\bigr).
\end{align}
The statement \eqref{eq:statement} is equivalent to the inequality
\begin{equation}
    (i+1)\alpha\beta\ge j+1.\label{eq:product}
\end{equation}
Indeed, if $(i+1)\alpha\ge k+1$ for some $k$, then $(i+1)\alpha\beta\ge (k+1)\beta$ since $\beta$ is order-preserving. Combining the latter inequality with $(k+1)\beta\ge j+1$ yields \eqref{eq:product}. Conversely, \eqref{eq:product} implies that \eqref{eq:statement} holds with $k:=(i+1)\alpha-1$.

The inequality \eqref{eq:product} is equivalent to $(i+1)\alpha\beta-1\ge j$, and the latter inequality is equivalent to $c_{ij}=1$ by the definition of the map $\alpha\mapsto M(\alpha)$. Hence $d_{ij}=1\iff c_{ij}=1$, and the matrices $M(\alpha)M(\beta)$ and $M(\alpha\beta)$ coincide.
\end{proof}

\begin{remark} The lower triangular Boolean matrices in the image of bijection $\alpha\mapsto M(\alpha)$ are in a natural one-to-one correspondence with the Young diagrams contained in the staircase shape $(n,n-1,\dots,1)$. The next illustration shows these matrices together with the corresponding Young diagrams (drawn in the French convention---rows are left-justified and counted from bottom to top) for $n=2$.

\def\boxsize{0.6cm}

\newcommand{\drawBounding}{%
  \draw[very thick,black] (0,0) rectangle ++(\boxsize,\boxsize);
  \draw[very thick,black] (\boxsize,0) rectangle ++(\boxsize,\boxsize);
  \draw[very thick,black] (0,\boxsize) rectangle ++(\boxsize,\boxsize);
}

\[
\begin{array}{cccccc}
&
\begin{pmatrix}0 & 0 \\ 0 & 0\end{pmatrix} &
\begin{pmatrix}0 & 0 \\ 1 & 0\end{pmatrix} &
\begin{pmatrix}0 & 0 \\ 1 & 1\end{pmatrix} &
\begin{pmatrix}1 & 0 \\ 1 & 0\end{pmatrix} &
\begin{pmatrix}1 & 0 \\ 1 & 1\end{pmatrix} \\[-.6em]
& & & & & \\
& \downarrow & \downarrow & \downarrow & \downarrow & \downarrow \\
& & & & & \\[-.6em]
& 
\begin{tikzpicture}
  \drawBounding;
\end{tikzpicture}
& 
\begin{tikzpicture}
   \fill[fill=lightgray] (0,0) rectangle ++(\boxsize,\boxsize);
   \drawBounding;
\end{tikzpicture}
& 
\begin{tikzpicture}
   \fill[fill=lightgray] (0,0) rectangle ++(\boxsize,\boxsize);
  \fill[fill=lightgray] (\boxsize,0) rectangle ++(\boxsize,\boxsize);
  \drawBounding;
\end{tikzpicture}
& 
\begin{tikzpicture}
   \fill[fill=lightgray] (0,0) rectangle ++(\boxsize,\boxsize);
  \fill[fill=lightgray] (0,\boxsize) rectangle ++(\boxsize,\boxsize);
  \drawBounding;
\end{tikzpicture}
& 
\begin{tikzpicture}
  \fill[fill=lightgray] (0,0) rectangle ++(\boxsize,\boxsize);
  \fill[fill=lightgray] (\boxsize,0) rectangle ++(\boxsize,\boxsize);
  \fill[fill=lightgray] (0,\boxsize) rectangle ++(\boxsize,\boxsize);
  \drawBounding;
\end{tikzpicture}
\end{array}
\]
\smallskip

\noindent The established bijection with $\mC^-_{n+1}$ shows that the number of Young diagrams contained in the staircase shape $(n,n-1,\dots,1)$ equals the $(n+1)$-st Catalan number. This fact is known; see \cite[Exercise 167]{Stanley}.
\end{remark}

\begin{cor}
\label{cor:C_(n+1)toT(n)}
The map $\alpha\mapsto PM(\alpha)P$ is a faithful representation of the semiring $\mC^-_{n+1}$ in the semiring $\mT_n$.
\end{cor}

\subsection*{Optimality} The faithful representation of Corollary \ref{cor:C_(n+1)toT(n)} is somewhat unexpected because of its `squeezing' feature: an object of a larger dimension embeds into one of smaller dimension. This raises the question: Can Catalan monoids or semirings be faithfully represented by Boolean triangular matrices of even smaller size? Our next results show that both Kl\'ima--Pol\'ak's embedding $\mC_n\hookrightarrow\mT_n$ and our embedding $\mC^-_{n+1}\hookrightarrow\mT_n$ already use triangular matrices of the minimum possible size.

\begin{theorem}
\label{thm:optimality}
\emph{(i)} The Catalan monoid $C_{n+2}$ does not embed into the monoid $T_n$.

\emph{(ii)} Neither of the Catalan semirings $\mC_{n+1}$ and $\mC^-_{n+2}$ embeds into the semiring $\mT_n$.
\end{theorem}

\begin{proof}
Our argument uses the fact that the identities
\begin{gather}
\label{eq:semigroup identity in Tn}x^n=x^{n+1},\\
\label{eq:semiring identity in Tn} x^{n-1}y^{n-1}=x^ny^{n-1}+x^{n-1}y^n
\end{gather}
hold in the semiring $\mT_n$ \cite[Example 2.4]{Vo25}.

Let $\alpha\in C_{n+2}$ be defined by
\[
i\alpha:=\begin{cases} i+1 &\text{if } i\le n+1,\\
n+2 &\text{if } i=n+2.
\end{cases}
\]
Then $1\alpha^n=n+1$ and $1\alpha^{n+1}=n+2$ whence $\alpha^n\ne\alpha^{n+1}$. Therefore, the identity \eqref{eq:semigroup identity in Tn} fails in the monoid $C_{n+2}$. As \eqref{eq:semigroup identity in Tn} holds in the monoid $T_n$ and identities are inherited by submonoids, $C_{n+2}$ does not embed in $T_n$ and the \ais{} $\mC^-_{n+2}$ whose multiplicative monoid is isomorphic to $C_{n+2}$ does not embed in $\mT_n$.

Let $\beta,\gamma\in C_{n+1}$ be defined by
\[
i\beta:=\begin{cases} i+1 &\text{if } i\le n,\\
n+1 &\text{if } i=n+1;
\end{cases}
\qquad
i\gamma:=\begin{cases} n &\text{if } i\le n,\\
n+1 &\text{if } i=n+1.
\end{cases}
\]
Then $1\beta^{n-1}\gamma^{n-1}=n\gamma^{n-1}=n$, whereas
\[
1\left(\beta^n\gamma^{n-1}+\beta^{n-1}\gamma^n\right)=\max\{1\beta^n\gamma^{n-1},1\beta^{n-1}\gamma^n\}=\max\{n+1,n\}=n+1.
\]
Therefore, the identity \eqref{eq:semiring identity in Tn} fails in the \ais{} $\mC_{n+1}$. As \eqref{eq:semiring identity in Tn} holds in $\mT_n$  and identities are inherited by subsemirings, $\mC_{n+1}$  does not embed in $\mT_n$.
\end{proof}

\begin{remark}
A \emph{divisor} of a monoid [semiring] is defined as a homomorphic image of one of its submonoids [subsemirings]. Since identities are inherited by divisors, the above proof of Theorem \ref{thm:optimality} shows that the results (i) and (ii) remain valid in the stronger form with embedding replaced by division.
\end{remark}

\subsection*{Complementarity} While the embedding $C^-_{n+1}\hookrightarrow T_n$ defined by $\alpha\mapsto PM(\alpha)P$ appears to be new, it is closely related to Kl\'ima--Pol\'ak's embedding and may be viewed, in a sense, as the complement of the embedding $C_{n+1}\hookrightarrow T_{n+1}$ defined by $\alpha\mapsto S(\alpha)$. Let us explain how this complementarity works.

It is easy to see that the lattice $(O_{n+1},\leqslant)$ is selfdual, the duality being the map $\alpha\mapsto\overline{\alpha}$ defined by
\[
i\overline{\alpha}:=n+2-(n+2-i)\alpha \ \text{ for each } \ i=1,2,\dots,n+1.
\]
The map $\alpha\mapsto\overline{\alpha}$ is known to be an automorphism of the monoid $O_{n+1}$, and the restriction of $\alpha\mapsto\overline{\alpha}$ to each of the Catalan monoids $C_{n+1}$ and $C^-_{n+1}$ is an isomorphism onto the other monoid. For $\alpha\in C_{n+1}$, `negate' the upper triangle of the $(n+1)\times(n+1)$-matrix $S(\alpha)$ by replacing all 1s in this triangle by 0s and all 0s with 1; the 0s below the main diagonal remain unchanged. Removing the first column and the last row from the resulting matrix yields an $n\times n$-matrix, which coincides with the matrix $PM(\overline{\alpha})P$.

The following illustration shows all steps of the process for $\alpha=1244\in C_4$. We restrict ourselves to this `proof by example' since a formal verification would merely require unfolding the definitions of all the maps involved.
\begin{center}
\begin{tikzpicture}[scale=1.4, every node/.style={minimum width=2cm, minimum height=2cm, align=center}]

\node (A) at (90:3)   {$\begin{pmatrix}1&0&0&0\\
0&1&0&0\\
0&0&1&1\\
0&0&0&1
\end{pmatrix}$};  
\node (B) at (30:3)   {$\begin{pmatrix}0&1&1&1\\
0&0&1&1\\
0&0&0&0\\
0&0&0&0
\end{pmatrix}$};  
\node (C) at (330:3)  {$\begin{pmatrix}1&1&1\\
0&1&1\\
0&0&0
\end{pmatrix}$};  
\node (D) at (270:3)  {$\begin{pmatrix}0&0&0\\
1&1&0\\
1&1&1
\end{pmatrix}$};  
\node (E) at (210:3)  {\begin{tikzpicture}[scale=0.5]
\useasboundingbox (1,2) rectangle (2.2,2.6); 
\foreach \x in {0.5,2.5} \node at (\x,0.5) {\tiny 1};
\foreach \x in {0.5,2.5} \node at (\x,1.5) {\tiny 2};
\foreach \x in {0.5,2.5} \node at (\x,2.5) {\tiny 3};
\foreach \x in {0.5,2.5} \node at (\x,3.5) {\tiny 4};
    \draw (0.75,0.5) edge[->] (2.25,0.5);
	\draw (0.75,1.5) edge[->] (2.25,0.5);
    \draw (0.75,2.5) edge[->] (2.25,2.5);
    \draw (0.75,3.5) edge[->] (2.25,3.5);
    \node at (1.5,-0.15)  {$\overline{\alpha}$};
\end{tikzpicture}};  
\node (F) at (150:3)  {\begin{tikzpicture}[scale=0.5]
\useasboundingbox (1,0) rectangle (2.2,2.6); 
\foreach \x in {0.5,2.5} \node at (\x,0.5) {\tiny 1};
\foreach \x in {0.5,2.5} \node at (\x,1.5) {\tiny 2};
\foreach \x in {0.5,2.5} \node at (\x,2.5) {\tiny 3};
\foreach \x in {0.5,2.5} \node at (\x,3.5) {\tiny 4};
    \draw (0.75,0.5) edge[->] (2.25,0.5);
	\draw (0.75,1.5) edge[->] (2.25,1.5);
    \draw (0.75,2.5) edge[->] (2.25,3.5);
    \draw (0.75,3.5) edge[->] (2.25,3.5);
\node at (1.5,4.15)   {$\alpha$};
\end{tikzpicture}};  

\draw[->, thick] (A) -- node[sloped, above, yshift=-4ex] {`negation'} (B);
\draw[->, thick] (B) -- node[align=center, right] {Removing the first\\ column and\\ the last row}(C);
\draw[<-, thick] (C) --  node[sloped, above, yshift=-4ex] {$PM(\overline{\alpha})P$} (D);
\draw[<-, thick] (D) --  node[sloped, above, yshift=-4ex] {$M(\overline{\alpha})$} (E);
\draw[<->, thick] (E) -- (F);
\draw[->, thick] (F) --  node[sloped, above, yshift=-4ex] {$S(\alpha)$} (A);

\end{tikzpicture}
\end{center}

\end{document}